OpenStudy (anonymous):

Solve the triangle. A = 51°, b = 11, c = 7 \

OpenStudy (anonymous):

Ever heard the Pythagorean Theorem? ^.^

OpenStudy (anonymous):

a2+b2=c2?

OpenStudy (anonymous):

well in this case, more like \[\large a^2=b^2+c^2...\]

OpenStudy (anonymous):

but this is no right triangle, so it doesn't really work out... :P

OpenStudy (anonymous):

that's where the LAW OF COSINES comes into play :D It's like the Pythagorean Theorem, but BETTER. It works for ANY triangle... ^.^ as long as you're either given all three sides or two sides and the angle between them, like here :P

OpenStudy (anonymous):

With just a reeaallly tiny addition to the formula, we get the law of Cosines ^.^\[\Large a^2 = b^2+c^2 \color{red}{-2bc \ \cos(A)}\]

OpenStudy (anonymous):

Now you can just solve for a :D

OpenStudy (anonymous):

12.8?

OpenStudy (anonymous):

@PeterPan